# An Analysis of the Impact of Gold Open Access Publications in Computer Science

### v1

@@ 1 Introduction
The world of scientific publishing has changed significantly in recent years. Two important changes are the increase in volume and the move to Open Access [12, 15]. The latter often requires the authors to pay a publishing fee, which has introduced new commercial opportunities and attracted new players into the publishing business. Some of these new publishers have been branded as ‘predatory’ [16], because they are viewed to be exploiting the Open Access model without following the best practices for academic publishing to ensure scientific quality. They often solicit and readily accept manuscripts by offering rapid review and publication, which can be problematic for maintaining the high scholarly standards needed to ensure lasting quality.
In this paper, we present an observational study of  1.4M computer science (CS) journal papers published between 2000 and 2023. The collection is based on the CS bibliography maintained on the DBLP website (dblp.org), and the citation data was obtained from Semantic Scholar (semanticscholar.org). We find that the number of journal publications per year in CS has increased from 20,000 in 2000 to 120,000 in 2023. While some of this increase can be accounted for by the ‘traditional’ journals from publishers such as Elsevier or IEEE, by 2023 one-third of publications now come from the new Open Access journals. This volume growth alone is concerning, but does it affect publication quality? To answer this question, we explore whether there are differences in quality between publications in new Open Access venues and those in traditional journals.
@@
  • We show that, on average, the impact of papers in Open Access journals is not equivalent to the impact of publications in traditional venues.
The paper proceeds as follows. In the next section, we provide a brief history of the emergence of Open Access publications in CS. In section 3, we review the main metrics used to quantify research impact. Section 4 provides details on the dataset used in our analysis. Section 5 presents the overall trends evident in the dataset, and section 6 discusses what can be learned about differences in citation impact from the data. The paper finishes with a summary and some conclusions in section 7.
@@ 2 Open Access Publication
There are at least two distinct forms of OA publishing [4]. With Gold OA the final published version of research articles is permanently and freely available and the authors will likely have paid an article processing charge (APC). With Green OA, a version of the article (but probably not the final version) is available in a repository such as arXiv, and no APC will have been paid.
OA benefits scientific progress and democratises research by making publications universally available. Because of this, it is research funding agencies increasingly require research results to be available as OA. However, despite the manifest benefits of OA, the initiative is an interesting example of the law of unintended consequences in action. Instead of readers paying for access to research papers, there has been a significant move towards researchers paying a significant and upfront charge for publication. This has created a large market for paid academic publication, with new opportunities to exploit researchers and academics who must survive in the increasingly competitive publish-or-perish academic world. Not surprisingly, this has attracted many new unscrupulous players into this market. Beall [4] argues that the Gold OA model has a fatal flaw: it motivates publishers to publish more and more but without the guardrails of scientific quality or research merit.
Indeed, the emergence of unethical practices in Gold OA has led to the widespread use of the terms ‘predatory journal’ and ‘predatory publishers’ to describe this phenomenon [7, 11]. A predatory publisher is a publisher whose main motivation is the financial gain that can be derived from a high-volume business model. The problem is exacerbated because the demarcation between legitimate and predatory publications is increasingly unclear. Publishers such as Hindawi, Frontiers and MDPI, which might have been branded predatory in the past, now host journals in their portfolios that are included in the main research indexing services such as PubMed and Scopus. Because of the emergence of this ‘grey area‘ between traditional and predatory journals, some publishers are referred to as ‘grey’ rather than ‘predatory’[16]. The landscape is further complicated because some traditional publishers have set up their own Gold OA journals to benefit from this increasingly lucrative market. IEEE Access is perhaps the most visible example of this, and for this reason, in our analysis, we distinguish between traditional IEEE journals and IEEE Access, grouping the latter with other grey publishers such as Hindawi, Frontiers, and MDPI.
@@ 3 Publication Metrics
It is worth emphasising that the best way to evaluate the quality of a research paper is to read it carefully [19]. At the same time, methods for quantifying the significance or impact of a paper do have an important and established role; for example, if we wish to read up on a new research area, then the most widely cited papers are a good place to start. The practice of ranking journals based on citation metrics is well established and relied upon by researchers, academic institutions, and research funding agencies as an important quality indicator. In this section, we present an overview of these metrics, focusing on metrics that might be impacted by an increase in publication volume and variability in publication quality.
@@ 3.2 $h$-Index
Another issue with the $h$-index (and with FWCI) is that it does account for differences in author counts. For example, with the $h$-index metric, every scholar in a multi-authored article receives full credit for its citation count, even though the contributions of all scholars are rarely equal. Accordingly, several proposed variants seek to normalise the $h$-index score based on author count [5].
@@ 3.4 Influential Citations
In the original Valenzuela and Etzioni work found that approximately 15% of citations were influential; the proportions we report in section 5 are somewhat lower than this. Crucially, this metric helps to differentiate between the Gold OA and other publications.
@@ 4 The Data
The dataset used in this study was generated using DBLP and Semantic Scholar (April-May, 2024). An initial set of papers was collected from DBLP by extracting all the journal articles listed, resulting in a collection of 1,896,440 unique articles.
@@
Next, we defined a set of Traditional publishers – IEEE (excluding IEEE Access), Elsevier, Springer, ACM) – and a set of Gold OA publishers – Hindawi, IEEE Access, Frontiers, MDPI – to distinguish between publishers adopting more conventional academic publication practices versus more recent paid, OA models. We identified all articles in our dataset published in these target journals. This resulted in a subset of 1,135,203 Traditional articles and 254,136 Gold OA articles; these articles are associated with 11,548,876 unique citations.
We use this 2000-2023 Traditional/Gold OA dataset in two ways: (i) to analyse overall trends in the CS publication landscape since 2000, and (ii) to analyse the citation impact of traditional versus OA publishers. Since traditional publishers have a longer history and, therefore, a greater opportunity to attract citations, for this impact analysis, we focus on a smaller subset of 850,369 articles published between 2015 and 2023; 620,807 Traditional and 229,562 Gold OA articles.
@@ 5 Overall CS Publication Trends
This section looks at the overall trends in the CS publication landscape since 2000, comparing Gold OA (Hindawi, MDPI, Frontiers, IEEE Access) and Traditional (IEEE, ACM, Elesvier, Springer) publishers. Altogether this comprises $\sim$ 1.4M papers. The main trends in this dataset are shown in Figures 1 and 2 and the following observations can be made.
@@
  • Paper Count: The first thing to note is the significant increase in publication volume (Figure 1). The output from the traditional publishers has grown steadily from about 20K papers in 2000 to over 80K p.a. in 2023. Significant output from Gold OA publishers emerged after 2008 and increased steeply in the five years up to 2020, reaching 40K papers p.a.
@@
  • Author Count: One of the more surprising trends in the dataset is shown in Figure 2(a). There has been a steady growth in the average number of authors per paper, growing from just over 2 authors per paper in 2000 to more than 4 authors per paper in 2023. This trend is evident in both the traditional and Gold OA collections. It is surprising because, unlike increases in references and citation counts, there is no obvious reason for it. This issue is discussed further in section 5.2.
@@
  • Reference Count: Figure 2(b) shows that the average number of references (cited papers) has been increasing steadily since 2000. This may be due to the move to online publication and the consequent relaxation of bibliographic limits in papers. However, it is noteworthy that this growth has continued since the early 2000s. While there is a sharp increase in the number of references in the Gold OA group from 2008 – coinciding with the rise of Gold OA papers in Figure 1 – which probably can be explained by more relaxed bibliographic limit, the reasons for the more stable growth among traditional journals appears less clear-cut.
@@
  • Citation Count: The plot in Figure 2(c) suggests that the two collections differ in terms of their mean citation count, but the picture is complicated by the fact that the citation count is lower for newer papers. Therefore, in Figure 2(d), we present a normalised version of citation count by dividing by the number of years since publication. Gold OA papers have historically attracted a lower average citation count than traditional venues, and while the situation improved somewhat around 2009, the difference remains.
@@
  • Influential Citation Count The difference between Gold OA and traditional publications is more apparent when we look at influential citation count in Figures 2(e) and 2(f); once again we present a normalised measure of influential citation count (2(f)) by using the number of years since publication as the normalisation factor. Papers published in traditional venues attract significantly more influential citations than those published in OA venues.
@@
In summary, there are several interesting and concerning things to note. The volume of publication has increased dramatically. The number of citations generated is increasing even more dramatically. Gold OA papers attract fewer citations than those in traditional venues, and the citations that they do attract are much less likely to be influential. Finally, the average number of authors per paper has almost doubled since 2000.
@@
Figure 1: The publication output of both traditional and Gold OA publishers has increased significantly since 2000.
@@ 5.1 Increase in Publication Volume
In Figure 1 we see that the number of publications has increased dramatically since 2000. To look at this in more detail, we look at the relative growth for different publishers since Gold OA became a significant factor from 2015 onward. In Figure 3 we show changes in publication volume by publisher for the two four-year periods before and after 2000.
The largest overall publishers are Elsevier and IEEE, in both periods, but the publishers with largest relative growth are Frontiers, MDPI and IEEE Access; since Frontiers is coming from a low base, its overall annual output remains small despite its 3.5x growth rate. It is worth noting that the current annual volume of journal papers from MDPI since 2020 is greater than that from Springer and the ACM combined, while IEEE Access now approaches the annual output of Springer and exceeds that of the ACM.
These volume changes have the potential to significantly reshape the CS landscape because the growth rate of the Gold OA publishers is more than twice that of traditional publishers. MDPI is now the number three producer of journal articles in CS, and even if it does not manage to sustain its current growth rate, even a reduced growth rate will likely be sufficient for it to overtake IEEE and perhaps Elsevier to become the largest producer of CS journal articles over the next four years.
@@ 5.2 The Increase in Author Count
We expect the average number of authors on papers to vary widely across disciplines, but there is no obvious reason for this authorship growth in Computer Science. It may be due to a move towards multi-disciplinary research or larger research groups, which may be characteristic of modern CS research.
As data-driven methods have significantly impacted the nature of science across many disciplines, it may be natural to find computer science authors as additional members of publication teams, with CS papers including scientists from other domains. This will continue as artificial intelligence and machine learning will likely play an increasingly important role in future research across all disciplines.
@@
To understand this change in authorship practice further, we have examined the data in more detail. Figure 4 presents the mean author count by publisher for the four years from 2020 to 2023. While three of the top four publishers are Gold OA, it is clear that the variation is not huge.
@@
Thus, we can conclude that there is a common trend toward larger authorship counts across all publishers. While some of this can be explained by ‘legitimate’ reasons, such as larger multi-disciplinary research teams, this is not always true. Therefore, it will be important to consider the impact of increasing authorship counts in the future, and since FWCI and $h$-index cannot do this, it will be important to consider variations in these metrics that can account for such trends – several of which have been proposed – or entirely new metrics.
@@ 6 Impact Differences
An important question in this paper relates to the changes in the CS publication landscape, particularly the increase in volume. Specifically, we are interested in whether publications and citations can be counted the same. We compare Gold OA and traditional publishers under two criteria: citation network centrality (introduced in Section 3.3) and influential citations (Section 3.4). The analysis is based on publications since 2015 because it is clear from Figure 1 that Gold OA publishers were not particularly active before 2015.

### v2

@@ 1 Introduction
The world of scientific publishing has changed significantly in recent years. Two important changes are the increase in volume and the move to Open Access [15, 18]. The latter often requires the authors to pay a publishing fee, which has introduced new commercial opportunities and attracted new players into the publishing business. Some of these new publishers have been branded as ‘predatory’ [19] because they are viewed to be exploiting the Open Access model without following the best practices for academic publishing to ensure scientific quality. They often solicit and readily accept manuscripts by offering rapid review and publication, which can be problematic for maintaining the high scholarly standards needed to ensure lasting quality.
In this paper, we present an observational study of  1.4M computer science (CS) journal papers published between 2000 and 2023. The collection is based on the CS bibliography maintained on the DBLP website (dblp.org), and the citation data was obtained from Semantic Scholar (semanticscholar.org). We find that the number of journal publications per year in CS has increased from 20,000 in 2000 to 120,000 in 2023. While some of this increase can be accounted for by the ‘traditional’ journals from publishers such as Elsevier or IEEE, by 2023 one-third of publications now come from the new Open Access journals. This volume growth alone is concerning, but does it affect publication quality? To answer this question, we explore whether there are differences in quality between publications in new Open Access venues and those in traditional journals. Our analysis accounts for the move to hybrid models that include OA among traditional publishers in order to distinguish between these and newer OA-first publications.
@@
  • We show that, on average, the impact of papers in New OA journals is not equivalent to the impact of publications in traditional venues.
  • Indeed, when traditional journals switch to a hybrid model (by accommodating OA papers) the impact of these OA papers holds up.
The paper proceeds as follows. In the next section, we provide a brief history of the emergence of Open Access publications in CS. In section 3, we review the main metrics used to quantify research impact. Section 4 provides details on the dataset used in our analysis. Section 5 presents the overall trends evident in the dataset, and section 6 discusses what can be learned about differences in citation impact from the data. The paper finishes with some recommendations for robust citation metrics in section 7 and a summary and some conclusions in section 8.
@@ 2 Open Access Publication
There are at least two distinct forms of OA publishing [4]. With Gold OA the final published version of research articles is permanently and freely available and the authors will likely have paid an article processing charge (APC). JMLR and JAIR would be unusual in this regard, these are Gold OA journals without an APC. With Green OA, a version of the article (but probably not the final version) is available in a repository such as arXiv, and no APC will have been paid.
OA benefits scientific progress and democratises research by making publications universally available. Because of this, research funding agencies increasingly require research results to be available as OA. However, despite the manifest benefits of OA, the initiative is an interesting example of the law of unintended consequences in action. Instead of readers paying for access to research papers, there has been a significant move towards researchers paying a significant and upfront charge for publication. This has created a large market for paid academic publication, with new opportunities to exploit researchers and academics who must survive in the increasingly competitive publish-or-perish academic world. Not surprisingly, this has attracted many new unscrupulous players into this market. Beall [4] argues that the Gold OA model has a fatal flaw: it motivates publishers to publish more and more but without the guardrails of scientific quality or research merit.
Indeed, the emergence of unethical practices in Gold OA has led to the widespread use of the terms ‘predatory journal’ and ‘predatory publishers’ to describe this phenomenon [9, 14]. A predatory publisher is a publisher whose main motivation is the financial gain that can be derived from a high-volume OA business model. The problem is exacerbated because the demarcation between legitimate and predatory publications is increasingly unclear. Publishers such as Hindawi, Frontiers and MDPI, which might have been branded predatory in the past, now host journals in their portfolios that are included in the main research indexing services such as PubMed and Scopus. Because of the emergence of this ‘grey area‘ between traditional and predatory journals, some publishers are referred to as ‘grey’ rather than ‘predatory’[19]. The landscape is further complicated because some traditional publishers have set up their own Gold OA journals to benefit from this increasingly lucrative market. IEEE Access is perhaps the most visible example of this, and for this reason, in our analysis, we distinguish between traditional IEEE journals and IEEE Access, grouping the latter with other grey publishers such as Hindawi, Frontiers, and MDPI. To further complicate things many traditional journals have adopted a hybrid model where papers are published Gold OA subject to an APC. Happily, our analysis suggests that there are no quality issues with these OA papers in hybrid journals.
@@ 3 Publication Metrics
It is worth emphasising that the best way to evaluate the quality of a research paper is to read it carefully [22]. At the same time, methods for quantifying the significance or impact of a paper do have an important and established role; for example, if we wish to read up on a new research area, then the most widely cited papers are a good place to start. The practice of ranking journals based on citation metrics is well established and relied upon by researchers, academic institutions, and research funding agencies as an important quality indicator. At the same time it must be emphasized that publication metrics are using citations as an indicator or proxy for research quality. There will be cases where good quality research is not picked up by these metrics. In this section, we present an overview of these metrics, focusing on metrics that might be impacted by an increase in publication volume and variability in publication quality.
@@ 3.2 $h$-Index
Another issue with the $h$-index (and with FWCI) is that it does not account for differences in author counts. For example, with the $h$-index metric, every scholar in a multi-authored article receives full credit for its citation count, even though the contributions of all scholars are rarely equal. Accordingly, several proposed variants seek to normalise the $h$-index score based on author count [6]. Two formulae for achieving this are as follows:
$h_{i}=\frac{h}{Avg_{A}};\quad h_{p}=\frac{h}{\sqrt{Avg_{A}}}$ (3)
$h_{i}$ is simply the $h$-index normalised by the average author count in the papers contributing to the index. $h_{p}$ is the pure $h$-index where the normalisation is the square root of the average author count – the objective being to dampen the impact of a small number of papers with very large author counts [6]. Given that papers with very long author lists are rare in Computer Science this square root may not be necessary; we return to this issue in section 7.
@@ 3.3 Network Centrality
In 2008, Bergstrom et al. introduced their Eigenfactor Metrics [5] as a concrete proposal for scoring journal impact using network centrality. Their Eigenfactor Score quantifies how important a journal is in a citation networks in terms of network centrality. Because this gives larger journals more influence, they propose the Article Influence Score which normalises the Eigenfactor Score by the number of articles published in the journal.
Another journal metric that uses network centrality is the SJR score[8]; it has the advantage that the SJR scores for most journals are freely available online at (scimagojr.com).
@@ 3.4 Influential Citations
In the original Valenzuela and Etzioni work found that approximately 15% of citations were influential; the proportions we report in section 5 are somewhat lower than this. Crucially, this metric helps to differentiate between the New OA and other publications.
@@ 4 The Data
The dataset used in this study was generated using DBLP and Semantic Scholar (April-May, 2024). An core set of papers was collected from DBLP by extracting all the journal articles listed, resulting in a collection of 1,896,440 unique articles. Thus DBLP defines the set of journal articles covered.
@@
Next, we defined a set of Traditional publishers – IEEE (excluding IEEE Access), Elsevier, Springer, ACM) – and a set of New OA publishers – Hindawi, IEEE Access, Frontiers, MDPI – to distinguish between publishers adopting more conventional academic publication practices versus more recent paid, OA models. We identified all articles in our dataset published in these target journals and further distinguished between OA articles from traditional publishers, which we refer to as Trad (OA), and non-OA articles from traditional publishers, which we refer to as Trad. This resulted in a subset of 755,890 Trad articles, and 379,313 Trad(OA) articles, and 254,136 New OA articles; these articles are associated with 11,548,876 unique citations.
We use this 2000-2023 Trad/Trad(OA)/New OA dataset in two ways: (i) to analyse overall trends in the CS publication landscape since 2000, and (ii) to analyse the citation impact of traditional versus (new) OA publishers. Since traditional publishers have a longer history and, therefore, a greater opportunity to attract citations, for this impact analysis, we focus on a smaller subset of 850,369 articles published between 2015 and 2023; 391,602 Trad, 229205 Trad (OA), and 229,562 New OA articles.
@@ 5 Overall CS Publication Trends
This section looks at the overall trends in the CS publication landscape since 2000, comparing New OA (Hindawi, MDPI, Frontiers, IEEE Access) and Traditional (IEEE, ACM, Elesvier, Springer) publishers. Altogether this comprises $\sim$ 1.4M papers. The main trends in this dataset are shown in Figures 1 and 2 and the following observations can be made.
@@
  • Paper Count: The first thing to note is the significant increase in publication volume (Figure 1). The output from the traditional publishers has grown steadily from about 20K papers in 2000 to over 80K p.a. in 2023. Significant output from New OA publishers emerged after 2008 and increased steeply in the five years up to 2020, reaching 40K papers p.a.
    There is a noticeable decline in the 2019 data for Traditional publishers; it is most obvious for Trad but also evident in Trad (OA). The 2020 numbers for Trad appear to be higher than expected, which may suggest that the previous year’s dip is due to a data issue with Semantic Scholar whereby some 2019 articles are incorrectly listed as 2020 articles. We do not believe that this issue has a material impact on our analysis of the results.
@@
  • Author Count: One of the more surprising trends in the dataset is shown in Figure 2(a). There has been a steady growth in the average number of authors per paper, growing from just over 2 authors per paper in 2000 to more than 4 authors per paper in 2023. This trend is evident in both the traditional and New OA collections. It is surprising because, unlike increases in references and citation counts, there is no obvious reason for it. This issue is discussed further in section 5.2.
@@
  • Reference Count: Figure 2(b) shows that the average number of references (cited papers) has been increasing steadily since 2000. This may be due to the move to online publication and the consequent relaxation of bibliographic limits in papers. However, it is noteworthy that this growth has continued since the early 2000s. While there is a sharp increase in the number of references in the New OA group from 2008 – coinciding with the rise of New OA papers in Figure 1 – which probably can be explained by more relaxed bibliographic limit, the reasons for the more stable growth among traditional journals appears less clear-cut. It is noteworthy too that the mean reference count for Trad (OA) papers exceeds that of Trad papers, just like the New (OA) papers.
@@
  • Citation Count: The plot in Figure 2(c) suggests that the two collections differ in terms of their mean citation count, but the picture is complicated by the fact that the citation count is lower for newer papers. Therefore, in Figure 2(d), we present a normalised version of citation count by dividing by the number of years since publication. New OA papers have historically attracted a lower average citation count than traditional venues, and while the situation improved somewhat around 2009, the difference remains. Moroever, the Trad (OA) papers have attracted a higher citation count than the non-OA Trad papers, consistent with the notion that OA papers are more readily discoverable than non-OA papers.
@@
  • Influential Citation Count The difference between New OA and traditional publications is more apparent when we look at influential citation count in Figures 2(e) and 2(f); once again we present a normalised measure of influential citation count (2(f)) by using the number of years since publication as the normalisation factor. Papers published in traditional venues attract significantly more influential citations than those published in OA venues. And, once again, Trad (OA) outperform Trad papers by this measure.
@@
In summary, there are several interesting and concerning things to note. Publication volume has increased dramatically. The number of citations generated is increasing even more dramatically. New OA papers attract fewer citations than those in traditional venues, but Trad OA papers attract more citations that New OA and Trad papers. Moroever, that the citations that the New OA papers do attract are much less likely to be influential, while the Trad (OA) citations are much more likely to be influential. This suggests an impact disparity that is due to the publisher type (grey versus traditional publishers) rather than the publishing model per se. Finally, the average number of authors per paper has almost doubled since 2000.
@@
Figure 1: The publication output of both traditional and New OA publishers has increased significantly since 2000.
@@ 5.1 Increase in Publication Volume
In Figure 1 we see that the number of publications has increased dramatically since 2000. To look at this in more detail, we look at the relative growth for different publishers since New OA became a significant factor from 2015 onward. In Figure 3 we show changes in publication volume by publisher for the two four-year periods before and after 2000.
The largest overall publishers are Elsevier and IEEE, in both periods, but the publishers with largest relative growth are Frontiers, MDPI and IEEE Access; since Frontiers is coming from a low base, its overall annual output remains small despite its 3.5$\times$ growth rate. It is worth noting that the current annual volume of journal papers from MDPI since 2020 is greater than that from Springer and the ACM combined, while IEEE Access now approaches the annual output of Springer and exceeds that of the ACM.
These volume changes have the potential to significantly reshape the CS landscape because the growth rate of the New OA publishers is more than twice that of traditional publishers. MDPI is now the number three producer of journal articles in CS, and even if it does not manage to sustain its current growth rate, even a reduced growth rate will likely be sufficient for it to overtake IEEE and perhaps Elsevier to become the largest producer of CS journal articles over the next four years.
@@ 5.2 The Increase in Author Count
We expect the average number of authors on papers to vary widely across disciplines, but there is no obvious reason for this authorship growth in Computer Science. It may be due to a move towards multi-disciplinary research or larger research groups, which may be characteristic of modern CS research. It may be due to increased collaboration between academic and industry partners.
As data-driven methods have significantly impacted the nature of science across many disciplines, it may be natural to find computer science authors as additional members of publication teams, with CS papers including scientists from other domains or from industry. This will continue as artificial intelligence and machine learning will likely play an increasingly important role in future research across all disciplines.
@@
To understand this change in authorship practice further, we have examined the data in more detail. Figure 4 presents the mean author count by publisher for the four years from 2020 to 2023. While three of the top four publishers are New OA, it is clear that the variation is not huge.
@@
Thus, we can conclude that there is a common trend toward larger authorship counts across all publishers. While some of this can be explained by ‘legitimate’ reasons, such as larger multi-disciplinary research teams, this is not always true. Therefore, it will be important to consider the impact of increasing authorship counts in the future, and since FWCI and $h$-index cannot do this, it will be important to consider variations in these metrics that can account for such trends; this issue is discussed in section 7.
@@ 6 Impact Differences
An important question in this paper relates to the changes in the CS publication landscape, particularly the increase in volume. Specifically, we are interested in whether publications and citations can be counted the same. We compare New OA and traditional publishers, on a publisher by publisher basis, under two criteria: citation network centrality (introduced in Section 3.3) and influential citations (Section 3.4). The analysis is based on publications since 2015 because it is clear from Figure 1 that New OA publishers were not particularly active before 2015.